# Differentiability of solutions for a degenerate fully nonlinear free transmission problem

### v1

@@
Abstract
We study a fully nonlinear free transmission problem in the presence of general degeneracy terms. Under minimal conditions on the degeneracy of the model, we establish the existence of viscosity solutions for the associated Dirichlet problem. Once the existence of solutions has been established, we focus on their regularity estimates. By imposing a Dini-continuity condition on the degeneracy laws involved in the model, we prove that viscosity solutions are locally differentiable.
Keywords: Degenerate fully nonlinear free transmission problems; existence of solutions; differentiability of solutions; regularity theory.
MSC(2020): 35B65; 35D40; 35A01.
1 Introduction
@@
We equip (1) with Dirichlet boundary data $g\in C(\partial\Omega)$ and prove the existence of viscosity solutions to the associated Dirichlet problem. Concerning the degeneracy rates $\sigma_{1}$ and $\sigma_{2}$, the existence of solutions only requires these functions to be monotone increasing and ordered. When it comes to regularity estimates, we impose a further condition. Namely, we require $\sigma_{2}$ to have an inverse $\sigma_{2}^{-1}$ which is itself a Dini modulus of continuity. Under such a condition, we prove the differentiability of the solutions to (1).
@@
The model in (1) amounts to a degenerate fully nonlinear free transmission problem. In particular, (1) describes a diffusion process degenerating as a modulus of continuity of the gradient. Meanwhile, the degeneracy depends on the sign of the solution.
@@
Discontinuous diffusions have been studied in the literature since the work of Mauro Picone, circa 1950; see [42]. The problem formulated in [42] finds its roots in the realm of elasticity theory. It attracted substantial attention from the mathematical community. Indeed, various authors have further developed the theory for that model, as well as proposed important variants. We refer the reader to [36, 34, 35, 47, 18, 19, 20, 45, 41, 48, 13, 26]; see also the monograph [14]. The main advances reported in these articles comprise the existence of solutions and their uniqueness and a priori estimates. After the establishment of a soundly based theory of the existence of solutions, the analysis of regularity properties took place. For example, we mention [33, 32, 5, 6].
More recently, a corpus of finer regularity results appeared in the context of transmission problems. In the interesting article [17], the authors examine the regularity of the solutions under minimal regularity conditions on the transmission interface. In that paper, the domain is split into two subdomains separated by a $C^{1,\alpha}$-regular interface. Under such an assumption, the authors first prove that solutions are locally of class $C^{0,{\rm Log-Lip}}$ in the domain. It stems from the basic properties of harmonic functions. Then they show that solutions are $C^{1,\alpha}$-regular up to the transmission interface. Their findings rely on a new stability result relating hypersurfaces of class $C^{1,\alpha}$ with flat ones.
In [46], the authors pursue the program launched in [17] in the context of fully nonlinear equations. Indeed, they develop a fairly complete theory of viscosity solutions for transmission problems governed by fully nonlinear operators. Their contributions include the existence and the uniqueness of solutions, an Aleksandrov-Bakelman-Pucci estimate, and regularity results. The main novelty in [46] regards the regularity of the solutions up to the transmission interface. As before, they prove the optimal regularity of the solutions matches the regularity of the interface in Hölder spaces.
@@
In recent years, the analysis of transmission problems started to account for diffusions with discontinuities across solution-dependent interfaces. That is, endogenously determined rather than prescribed a priori. The first analysis of a free transmission problems appeared recently in [1]. See also [40, 39].
In [21], the authors propose a variational free transmission problem modelled after the $p$-Laplace operator. They prove the existence of a minimiser $u$ and obtain Hölder-continuity estimates. In addition, they examine the associated free boundary. Among their findings, we highlight a free boundary condition and the almost everywhere $C^{1,\alpha}$-regularity with respect to a $p$-harmonic measure. Finally, the authors verify the support of $\Delta_{p}u^{+}$ is of $\sigma$-finite $(d-1)$-dimensional Hausdorff measure.
@@
Free transmission problems driven by uniformly elliptic fully nonlinear operators have been addressed in [44] and [43]. In [44], the authors prove the existence of $L^{p}$-viscosity solutions and $L^{p}$-strong solutions for the problem
$\begin{split}F_{1}(D^{2}u)=f&\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega\cap\left\{u>0\right\}\\ F_{2}(D^{2}u)=f&\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega\cap\left\{u<0\right\},\end{split}$ (2)
where $F_{i}:S(d)\to\mathbb{R}$ are uniformly elliptic operators. The regularity theory for (2) is the subject of [43]. In that paper, the authors prove that solutions to (2) are in $W^{2,{\rm BMO}}(\Omega)$, with estimates. Under a smallness assumption on the density of the negative phase, they obtain $C^{1,1}$-regularity estimates.
In the recent paper [31], the authors launch the analysis of fully nonlinear parabolic free transmission problems. Their main result concerns the regularity of the free boundary, as they verify that flat free boundaries are smooth.
Degenerate fully nonlinear free transmission problems are the subject of [25]. The authors consider
@@
The model in (3) is motivated by the study of fully nonlinear equations degenerating as a power of the gradient. See [7, 8, 9, 10, 11, 23], to name just a few. The work-horse of this theory is the equation
$|Du|^{\theta}F(D^{2}u)=f\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega,$ (4)
where $\theta\in(-1,\infty)$ is a fixed constant, $F$ is a uniformly elliptic operator, and $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$. It is well-known that solutions to (4) are locally $C^{1,\alpha^{*}}$-regular, where $\alpha^{\ast}\in(0,1)$ satisfies
$\alpha^{*}<\min\left(\alpha_{0},\frac{1}{1+\theta}\right),$
and $\alpha_{0}\in(0,1)$ is the universal exponent in the Krylov-Safonov theory available for $F=0$. See [27, 12, 3]. Further developments include regularity estimates for equations holding in the regions where the gradient is large and the case of variable exponents; see also [28, 37, 15]. For optimal boundary regularity estimates for the solutions to (4), see the developments reported in [4].
@@
In the present paper, we extend the approach in [25] to a degenerate transmission problem governed by equations as in (5). We start with a result on the existence of solutions. Let $g\in C(\partial\Omega)$ and consider the Dirichlet problem
[…]
Our first main result provides conditions on the degeneracy laws $\sigma_{1}$ and $\sigma_{2}$ ensuring the existence of viscosity solutions to (6). This is the content of the next theorem.
Theorem 1 (Existence of viscosity solutions).
Let $\Omega\subset\mathbb{R}^{d}$ be a domain satisfying a uniform exterior sphere condition. Suppose $F$ is a $(\lambda,\Lambda)$-uniformly elliptic operator, $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$, and $g\in C(\partial\Omega)$. Suppose further that $\sigma_{i}\in C(\mathbb{R}_{+})$ are monotone increasing, with $\sigma_{2}(t)\leq\sigma_{1}(t)$ for every $t>0$, and $\sigma_{i}(t)\to 0$ as $t\to 0$. Then there exists a viscosity solution $u\in C(\Omega)$ to (6). Furthermore, $u$ is a viscosity sub-solution to
$\min\left\{\sigma_{1}(|Du|)F(D^{2}u),\sigma_{2}(|Du|)F(D^{2}u)\right\}=\left\|f\right\|_{L^{\infty}(\Omega)}\hskip 7.22743pt\mbox{in}\hskip 7.22743pt\Omega$ (7)
and a viscosity super-solution to
$\max\left\{\sigma_{1}(|Du|)F(D^{2}u),\sigma_{2}(|Du|)F(D^{2}u)\right\}=-\left\|f\right\|_{L^{\infty}(\Omega)}\hskip 7.22743pt\mbox{in}\hskip 7.22743pt\Omega.$ (8)
The proof of Theorem 1 relies on an approximate problem. Indeed, to the best of our knowledge, the comparison principle is not available for viscosity solutions of (6). This is expected since the dependence of the equation in (6) is not monotone. Therefore, there is no a priori reason to ensure properness of the operator. We design an approximate problem depending on a functional parameter $v\in C(\overline{\Omega})$ and a perturbation parameter $\varepsilon>0$. For such a problem, the comparison principle is available; hence, we construct sub and super-solutions agreeing with $g$ on $\partial\Omega$. An application of Perron’s method yields the existence of a viscosity solution for $v$ and $\varepsilon$ fixed. Schauder’s fixed point theorem eliminates the functional parameter, whereas the stability of viscosity solutions allows us to take the limit $\varepsilon\to 0$, ultimately establishing the Theorem 1.
We stress two interesting consequences of our arguments. First, a simple adjustment of our methods provides the existence of viscosity solutions to (5), completing the program launched in [2]. Also, by modifying the sub and super-solutions used in the proof of Theorem 1, we learn the solutions whose existence follows from the theorem are indeed Lipschitz continuous. We detail these remarks further in the paper.
Once the existence of solutions has been understood, we shift our focus to the regularity of the solutions. Here, we suppose that $\sigma_{2}\leq\sigma_{1}$ and impose a Dini-continuity condition on $\sigma_{2}^{-1}$. Under those conditions, we prove that solutions to (1) are locally differentiable. Without loss of generality, we set $\Omega\equiv B_{1}$ when dealing with regularity estimates. Our second main theorem reads as follows.
@@
Let $u\in C(B_{1})$ be a viscosity solution to (1). Suppose $F$ is a $(\lambda,\Lambda)$-uniformly elliptic operator and $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$. Suppose also $\sigma_{i}$ are monotone increasing, with $\sigma_{2}(t)\leq\sigma_{1}(t)$ for every $t>0$. Suppose further that $\sigma_{2}^{-1}$ is a Dini-continuous modulus of continuity. Then $u\in C^{1}_{{\rm loc}}(B_{1})$. Moreover, there exists a modulus of continuity $\omega:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$ depending only upon the dimension, ellipticity constants, $\sigma_{i}$, $\|u\|_{L^{\infty}(B_{1})}$, and $\|f\|_{L^{\infty}(B_{1})}$ such that
@@
The proof of Theorem 2 relies on several ingredients. First, we notice that solutions to (1) satisfy two viscosity inequalities in the entire domain $B_{1}$. Then we resort to standard approximation methods to construct hyperplanes locally comparable with the solution, except for a prescribed error. Finally, the summability due to the Dini continuity of $\sigma_{2}^{-1}$ ensures the convergence of such hyperplanes and concludes the proof. A by-product of our argument is an explicit modulus of continuity for $Du$.
We notice that it suffices to require $\sigma_{2}^{-1}$ to be a Dini-continuous modulus of continuity, with no similar requirement on $\sigma_{1}$. This is because we resort to a characterisation of Dini-continuity based on the summability of a series governed by $\sigma_{2}^{-1}$. Therefore, the assumption that $\sigma_{2}\leq\sigma_{1}$ builds upon the Dini-continuity of $\sigma_{2}^{-1}$ to ensure that $\sigma_{1}^{-1}$ is also Dini-continuous.
The remainder of this paper is organised as follows. Section 2.1 details our main assumptions, whereas Section 2.2 gathers definitions and former results used in the paper. The existence of solutions is the subject of Section 3, where we detail the proof of Theorem 1. In Section 4 we put forward the proof of Theorem 2, establishing the optimal regularity of the solutions to (1).
2 Preliminaries
@@
2.1 Main assumptions
When studying the existence of solutions for the Dirichlet problem associated with (1), we impose a geometric condition on the boundary of $\Omega$.
Definition 1 (Uniform exterior sphere condition).
Let $\Omega\subset\mathbb{R}^{d}$ be a domain. We say that $\Omega$ satisfies a uniform exterior sphere condition if for any $x\in\partial\Omega$ there exists a ball $B\subset\mathbb{R}^{d}\setminus\Omega$ of radius $r>0$, independent of $x$, such that $x\in\partial B$.
Our first assumption concerns the geometry of $\partial\Omega$.
Assumption 1 (Uniform exterior sphere condition).
We suppose the domain $\Omega\subset\mathbb{R}^{d}$ satisfies a uniform exterior sphere condition.
We proceed by requiring the fully nonlinear operator $F$ to be uniformly elliptic.
@@ 2.1 Main assumptions
The existence of solutions relies on a monotonicity condition on the degeneracy rates. This is the content of our next assumption.
Assumption 3 (Monotonicity of degeneracy rates).
@@
for $i=1,2$. In addition,
$\sigma_{1}(t)\geq\sigma_{2}(t)$
for every $t\in[0,1]$. Furthermore, we assume without loss of generality that
$\sigma_{1}(1)\geq\sigma_{2}(1)\geq 1.$
When it comes to regularity estimates, additional conditions on $\sigma_{2}$ are necessary. Namely, one requires this degeneracy law to have an inverse which is a Dini-continuous modulus of continuity.
Assumption 4 (Dini continuity).
The degeneracy rate $\sigma_{2}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ has an inverse $\sigma_{2}^{-1}$ that is a Dini-continuous moduli of continuity.
A typical example of degeneracy law satisfying Assumptions 3 and 4 is
$\sigma_{1}(t):=t^{p_{1}}\hskip 21.68121pt\mbox{and}\hskip 21.68121pt\sigma_{2}(t):=t^{p_{2}},$
where $1\leq p_{2}<p_{1}$. They satisfy Assumption 3. Also, $\sigma_{i}^{-1}(t)=t^{1/p_{i}}$, which is Hölder continuous with exponent $1/p_{i}$. In what follows, we gather auxiliary results used in the manuscript.
2.2 Elementary notions and auxiliary results
We start with the notion of degenerate ellipticity. More general than the requirements in Assumption 2, this notion unlocks an important variant of the maximum principle, which we recall later.
Definition 2 (Degenerate ellipticity).
We say the operator $G:S(d)\times\mathbb{R}^{d}\times\mathbb{R}\times\Omega\to\mathbb{R}$ is degenerate elliptic if, for every $(p,r,x)\in\mathbb{R}^{d}\times\mathbb{R}\times\Omega$ and $M,\,N\in S(d)$ with $M\leq N$, we have
$G(M,p,r,x)\leq G(N,p,r,x).$
Degenerate ellipticity can be regarded as an order-preserving property of a fully nonlinear operator. We note it follows from uniform ellipticity. For completeness, we recall the notion of viscosity solution.
@@ 2.2 Elementary notions and auxiliary results
$G(D^{2}u,Du,u,x)=0\hskip 21.68121pt\mbox{in}\hskip 21.68121pt\Omega$ (9)
@@
$\mathcal{A}_{\lambda,\Lambda}:=\left\{A\in S(d)\;|\;\lambda|\xi|^{2}\leq A\xi\cdot\xi\leq\Lambda|\xi|^{2}\hskip 7.22743pt\mbox{for every}\hskip 7.22743pt\xi\in\mathbb{R}^{d}\right\}.$
@@
$\mathcal{M}^{-}_{\lambda,\Lambda}(M):=\inf_{A\in\mathcal{A}_{\lambda,\Lambda}}\,{\rm Tr}(AM).$
Also, define $\mathcal{M}^{+}_{\lambda,\Lambda}(M):=-\mathcal{M}^{-}_{\lambda,\Lambda}(-M)$.
For properties of the extremal operators, we refer the reader to [16, Lemma 2.10]. Among other things, the extremal operators are useful in defining uniform ellipticity. Indeed, an operator $F:S(d)\to\mathbb{R}$ satisfies Assumption 2 if, for any $M,\,N\in S(d)$, we have
@@
In the study of interior regularity of (1), we resort to a sequential argument (usual in the setting of approximation methods). A fundamental ingredient in this realm is compactness. We continue by recalling auxiliary results used to produce compactness in our context.
@@
For a proof of Proposition 1, we refer the reader to [22, Theorem 3.2]. We continue with a variant of a regularity result for equations holding only where the gradient is large. See [29]; see also [38].

### v2

@@
Abstract.
We study a fully nonlinear free transmission problem in the presence of general degeneracy laws. Under minimal conditions on the degeneracy of the model, we establish the differentiability of viscosity solutions.
Key words and phrases:
Degenerate fully nonlinear free transmission problems; differentiability of solutions; Dini continuity.
2020 Mathematics Subject Classification:
35B65.
1. Introduction
@@
Concerning the degeneracy rates $\sigma_{1}$ and $\sigma_{2}$, we require these functions to be monotone. We also suppose $\sigma_{1}$ and $\sigma_{2}$ to have inverses $\sigma_{1}^{-1}$ and $\sigma_{2}^{-1}$, which are themselves Dini moduli of continuity. Under such a condition, we prove the differentiability of the solutions to (1).
@@
The model in (1) amounts to a degenerate fully nonlinear free boundary problem. In particular, (1) describes a diffusion process degenerating as a modulus of continuity of the gradient. Meanwhile, the degeneracy depends on the sign of the solution.
@@
Discontinuous diffusions have been studied in the literature since the work of Mauro Picone, circa 1950; see [19]. The problem formulated in [19] finds its roots in the realm of elasticity theory. We refer the reader to [5]. After the establishment of a soundly based theory of the existence of solutions, the analysis of regularity properties took place. For example, we mention [17, 16].
More recently, a corpus of finer regularity results appeared in the context of transmission problems. Here, the authors examine the regularity of the solutions under minimal regularity conditions on the transmission interface; see [7, 22],
@@
In recent years, the analysis of transmission problems started to account for diffusions with discontinuities across solution-dependent interfaces. The analysis of a variational, uniformly elliptic free transmission problem appeared recently in [1]. See [8] for a variational free transmission problem modelled after the $p$-Laplace operator.
@@
Free transmission problems driven by uniformly elliptic fully nonlinear operators have been addressed in [21], [20], and [15]. Degenerate fully nonlinear free transmission problems are the subject of [11]. The authors consider
@@
The model in (2) is motivated by the study of fully nonlinear equations degenerating as a power of the gradient. The workhorse of the theory is the equation
(3) $|Du|^{\theta}F(D^{2}u)=f\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega,$
where $\theta\in(-1,\infty)$ is a fixed constant, $F$ is a uniformly elliptic operator, and $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$. It is well-known that solutions to (3) are locally $C^{1,\alpha}$-regular; see [12, 4, 3].
@@
The connection of (3) and (2) is subtle. The seminal contribution connecting these models appeared in [6]. In that paper, the authors examine a state-dependent degeneracy rate $\theta:\Omega\to\mathbb{R}$, giving rise to
(5) $|Du|^{\theta(x)}F(D^{2}u)=f\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega.$
In [6], the authors establish the optimal regularity of the solutions in $C^{1,\alpha}$ spaces. A delicate argument yields optimal regularity estimates not depending on the modulus of continuity of $\theta(\cdot)$. This fine analysis frames degenerate free boundary problems as a limit case of a model driven by continuous variable exponents. Compare (2) with (5), where $\theta(x)$ is an approximation of the exponential rule switching from $\theta_{1}$ to $\theta_{2}$.
Motivated by the ideas launched in [6], we extend the approach in [11] to (1). Without loss of generality, we set $\Omega\equiv B_{1}$ and state our main result as follows.
@@
Let $u\in C(B_{1})$ be a viscosity solution to (1). Suppose $F$ is a $(\lambda,\Lambda)$-uniformly elliptic operator and $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$. Suppose also $\sigma_{i}$ are monotone increasing. Suppose further that $\sigma_{i}^{-1}$ is a Dini-continuous modulus of continuity. Then $u\in C^{1}_{{\rm loc}}(B_{1})$. Moreover, there exists a modulus of continuity $\omega:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$ depending only upon the dimension, ellipticity constants, $\sigma_{i}$, $\|u\|_{L^{\infty}(B_{1})}$, and $\|f\|_{L^{\infty}(B_{1})}$ such that
@@
The proof of Theorem 1 relies on several ingredients. First, we notice that solutions to (1) satisfy two viscosity inequalities in the entire domain $B_{1}$. Then we resort to standard approximation methods to construct hyperplanes locally comparable with the solution, except for a prescribed error. Finally, the summability due to the Dini continuity of $\sigma_{i}^{-1}$ ensures the convergence of such hyperplanes and concludes the proof. A by-product of our argument is an explicit modulus of continuity for $Du$.
The remainder of this paper is organised as follows. Section 2 details our main assumptions and gathers definitions and former results used in the paper. The proof of Theorem 1 is the subject of Section 3.
2. Preliminaries
@@
We continue with an assumption on the degeneracy laws.
Assumption 2 (Degeneracy rates).
@@
for $i=1,2$. In addition, they have inverses $\sigma_{1}^{-1}$ and $\sigma_{2}^{-1}$ that are Dini-continuous moduli of continuity.
A typical example of degeneracy laws satisfying Assumptions 2 is
$\sigma_{1}(t)\coloneqq t^{p_{1}}\hskip 21.68121pt\mbox{and}\hskip 21.68121pt\sigma_{2}(t)\coloneqq t^{p_{2}},$
Also, $\sigma_{i}^{-1}(t)=t^{1/p_{i}}$, which is Hölder continuous with exponent $1/p_{i}$. In what follows, we gather auxiliary results used in the paper. We continue with the notion of viscosity solutions.
@@
(6) $G(D^{2}u,Du,u,x)=0\hskip 21.68121pt\mbox{in}\hskip 21.68121pt\Omega$
@@
$\mathcal{A}_{\lambda,\Lambda}\coloneqq\left\{A\in S(d)\;|\;\lambda|\xi|^{2}\leq A\xi\cdot\xi\leq\Lambda|\xi|^{2}\hskip 7.22743pt\mbox{for every}\hskip 7.22743pt\xi\in\mathbb{R}^{d}\right\}.$
@@
$\mathcal{M}^{-}_{\lambda,\Lambda}(M)\coloneqq\inf_{A\in\mathcal{A}_{\lambda,\Lambda}}\,{\rm Tr}(AM).$
Also, define $\mathcal{M}^{+}_{\lambda,\Lambda}(M)\coloneqq-\mathcal{M}^{-}_{\lambda,\Lambda}(-M)$.
Among other things, the extremal operators are useful in defining uniform ellipticity. Indeed, an operator $F:S(d)\to\mathbb{R}$ satisfies Assumption 1 if, for any $M,\,N\in S(d)$, we have
@@
We continue by recalling auxiliary results used to produce compactness in our context.
@@
For a proof of Proposition 1, we refer the reader to [9, Theorem 3.2]. We continue with a variant of a regularity result for equations holding only where the gradient is large. See [13]; see also [18].